Good afternoon sir/madam
Please assist me with this problem:
If 2 workers take 4 hours to process 100 items, how long will 5 workers take to process the same quantity?

Thank you very much for your assistance.
Rgds
Deals

Found 2 solutions by stanbon, Edwin McCravy:
Answer by stanbon(75887)

If 2 workers take 4 hours to process 100 items, how long will 5 workers take to process the same quantity?
----
#of workers and hours are indirectly related.
---
General Equation:
time = k/n where n is number of workers.
---
Solve for "k" using "2 workers take 4 hours":
4 = k/2
k = 8
-----
Specific Equation for this Problem:
time = 8/n
----
Now solve time if n = 5 workers.
t = 8/5
---
time = (8/5)hrs = 1 3/5 hrs = 1hr 36 minutes
==============================================
Cheers,
Stan H.
==========

Answer by Edwin McCravy(20055)

There are several ways you could think it out. Here are two:

If 2 workers take 4 hours to process them, then 1 worker would take twice as 
long, or 8 hours.  Therefore it would take 5 workers only one-fifth as long,
and one-fifth of 8 is 8/5 or 1 3/5 hours or 1 hour and 36 minutes.

Or

If 2 workers take 4 hours to process them, then 10 workers could do that
in one-fifth the time or 4/5 of an hour, and only 5 workers would take twice
that long, or 8/5 or 1 3/5 hours or 1 hour and 36 minutes.

Edwin